# On the Existence and Behaviour of Secondary Attention Sinks

### v1

@@ 1 Introduction
Attention sinks were first identified by first-attn-sink, where the BOS token was observed to receive anomalously high attention weights. This phenomenon has since shown broad practical implications, including LLM quantization (prefixing-attn-sink; liu2024intactkv), KV-cache optimization (rkv; cai2024pyramidkv), efficient LLM serving (first-attn-sink), and model enhancement (calibration-sink).
Many recent studies investigated the formation and functional role of the BOS sink. cancedda2024spectral analyzes attention sinks from a spectral subspace perspective, while When-attention-sink-emerges interprets them as a form of positional bias that mitigates over-mixing. Building on this line of work, two-sides-of-the-same-coin further examines their role in the depth-wise organization of the model, proposing that attention sinks serve as a mechanism for information compression along the depth dimension.
massive-activation and calibration-sink show that attention sinks are not limited to the BOS token; instead, multiple tokens can function as attention sinks. What-are-you-sinking further analyzes this phenomenon from a geometric perspective, showing that the emergence of multiple attention sinks is closely tied to the model’s positional embedding scheme. In this view, attention sinks act as reference points in a high-dimensional representation space, enabling the model to establish a stable internal coordinate system. However, the multiple sinks identified in previous work are fundamentally the same as the BOS sink: they emerge at the same layers and persist throughout the network. In contrast, we identify a new type of sink that differs in both its layer of emergence and its lifetime.
@@ 2 Background and Preliminary
Position embeddings $P$ provide tokens with positional information in the attention mechanism. Common $P$ include absolute positional embeddings (attention-is-all-you-need), rotary positional embeddings (RoPE) (rope), and NTK [^1] NTK stands for Neural Tangent Kernel, this refers to changing the base frequency in RoPE based on the network capability, allowing improved performance on long-context sequences.-aware scaled RoPE (yarn; effective-long). RoPE is used in models such as LLaMA llama2 and Mistral mistral7b, while NTK-aware scaled RoPE is adopted by Qwen team2024qwen2 and CodeLlama Codellama. Compared to RoPE, NTK-aware scaled RoPE increases the rotary base resulting in less rotation with position effective-long.
@@
Attention sink refers to the phenomenon where semantically uninformative tokens are assigned disproportionately high attention weights across diverse model architectures and scales, resulting in vertical patterns in the attention weight matrix (first-attn-sink).
Recent work What-are-you-sinking suggests that reducing the rotary frequency weakens the dot-product advantage of the first token. As a result, standard RoPE often produces a strong attention sink at the BOS token, whereas NTK-aware scaled RoPE can give rise to additional attention sinks at other positions in the sequence. In this work, we evaluate 11 model families that use distinct rotary base frequencies (10K-500K). Our findings on QwQ and the Qwen2/2.5/3 model families aligns with What-are-you-sinking. However, we observe that CodeLlama and several other model families, despite using very large rotary bases, do not exhibit such secondary sinks. The underlying cause of this discrepancy remains an open question.
@@ 3 Properties of Secondary Sinks
We perform extensive empirical experiments to study the properties of Secondary Sinks by comparing them with the BOS sink. To observe the Secondary Sinks, we generate reasoning traces from LLMs (DeepSeek-14B, DeepSeek-32B, Qwen-3-32B) on datasets including AIME24 (AIME2024), Math (math-500), and then threshold the attention weights to locate sink tokens across different layers. We find that these attention sinks consistently exhibit at least an order of magnitude higher $\ell_{2}$-norms than average token [^2] Tokens excluding attention sinks. and align along the similar the BOS sink direction.
@@ 4 Causal Formation Of Secondary Sinks
These results indicate that the sink direction is encoded within the MLP layers, and that future secondary sink tokens contain the information needed to recover this direction. Motivated by the linear hypothesis (park2023linear), we hypothesize that these future secondary sinks share a common key can retrieve the sink direction.
@@ 5 Secondary Sinks Effect on Attention
We follow (two-sides-of-the-same-coin) to define the attention sink-score for token position $k$ at layer $l$ as:

### v2

# On the Existence and Behavior of Secondary Attention Sinks

@@ 1 Introduction
Attention sinks were first identified by Xiao et al. (2023), where the BOS token was observed to receive anomalously high attention weights. This phenomenon has since shown broad practical implications, including LLM quantization (Son et al., 2024; Liu et al., 2024), KV-cache optimization (Cai et al., 2025; 2024), efficient LLM serving (Xiao et al., 2023), and model enhancement (Yu et al., 2024).
Many recent studies investigated the formation and functional role of the BOS sink. Cancedda (2024) analyzes attention sinks from a spectral subspace perspective, while Gu et al. (2024) interprets them as a form of positional bias that mitigates over-mixing. Building on this line of work, Queipo-de-Llano et al. (2025) further examines their role in the depth-wise organization of the model, proposing that attention sinks serve as a mechanism for information compression along the depth dimension.
Sun et al. (2024) and Yu et al. (2024) show that attention sinks are not limited to the BOS token; instead, multiple tokens can function as attention sinks. Ruscio et al. (2025) further analyzes this phenomenon from a geometric perspective, showing that the emergence of multiple attention sinks is closely tied to the model’s positional embedding scheme. In this view, attention sinks act as reference points in a high-dimensional representation space, enabling the model to establish a stable internal coordinate system. However, the multiple sinks identified in previous work are fundamentally the same as the BOS sink: they emerge at the same layers and persist throughout the network. In contrast, we identify a new type of sink that differs in both its layer of emergence and its lifetime.
@@ 2 Background and Preliminary
Position embeddings $P$ provide tokens with positional information in the attention mechanism. Common $P$ include absolute positional embeddings (Vaswani et al., 2017), rotary positional embeddings (RoPE) (Su et al., 2024), and NTK [^1] NTK stands for Neural Tangent Kernel, this refers to changing the base frequency in RoPE based on the network capability, allowing improved performance on long-context sequences.-aware scaled RoPE (Peng et al., 2023; Xiong et al., 2024). RoPE is used in models such as LLaMA (Touvron et al., 2023) and Mistral (Jiang et al., 2023), while NTK-aware scaled RoPE is adopted by Qwen (Team and others, 2024) and CodeLlama (Roziere et al., 2023). Compared to RoPE, NTK-aware scaled RoPE increases the rotary base resulting in less rotation with position (Xiong et al., 2024).
@@
Attention sink refers to the phenomenon where semantically uninformative tokens are assigned disproportionately high attention weights across diverse model architectures and scales, resulting in vertical patterns in the attention weight matrix (Xiao et al., 2023).
Recent work Ruscio et al. (2025) suggests that reducing the rotary frequency weakens the dot-product advantage of the first token. As a result, standard RoPE often produces a strong attention sink at the BOS token, whereas NTK-aware scaled RoPE can give rise to additional attention sinks at other positions in the sequence. In this work, we evaluate 11 model families that use distinct rotary base frequencies (10K-500K). Our findings on QwQ and the Qwen2/2.5/3 model families aligns with Ruscio et al. (2025). However, we observe that CodeLlama and several other model families, despite using very large rotary bases, do not exhibit such secondary sinks. The underlying cause of this discrepancy remains an open question.
@@ 3 Properties of Secondary Sinks
We perform extensive empirical experiments to study the properties of Secondary Sinks by comparing them with the BOS sink. To observe the Secondary Sinks, we generate reasoning traces from LLMs (DeepSeek-14B, DeepSeek-32B, Qwen-3-32B) on datasets including AIME24 (MAA, 2024), Math (Hendrycks et al., 2021), and then threshold the attention weights to locate sink tokens across different layers. We find that these attention sinks consistently exhibit at least an order of magnitude higher $\ell_{2}$-norms than average token [^2] Tokens excluding attention sinks. and align along the similar the BOS sink direction.
@@ 4 Causal Formation Of Secondary Sinks
These results indicate that the sink direction is encoded within the MLP layers, and that future secondary sink tokens contain the information needed to recover this direction. Motivated by the linear hypothesis (Park et al., 2023), we hypothesize that these future secondary sinks share a common key can retrieve the sink direction.
@@ 5 Secondary Sinks Effect on Attention
We follow Queipo-de-Llano et al. (2025) to define the attention sink-score for token position $k$ at layer $l$ as: